OpenStudy (anonymous):

Determine a quadratic function f(x)= ax^2 + bx + c if its graph passes through the point (2,19) and it has a horizontal tangent at (-1, -8)

hero (hero):

After substituting both points into the formula you get the following: 19 = -4a + 2b + c -8 = -a -b + c Solving for c, we get: c = a + b - 8 If we substitute c back into the first equation above, we get: 19 = -4a + 2b + (a + b - 8) 19 = -3a +3b - 8 27 = -3a + 3b 27+3a = 3b (27+3a)/3 = b 9 + a = b When we take the derivative and set f'(x) = 0, we end up with: x = -b/2a Now we substitute -1 for x, since it is the x vertex of the quad function along with b = (9 + a): -1 = -(9+a)/2a -1 = (-9-a)/2a -2a = -9-a 9+a = 2a 9 = a So: 9 = a 18 = b Now we do actually have enough information to find c: Going back to the first formula: -19 = -4(9) + 2(18) + c 19 = -36 + 36 + c 19 = c Therefore, the original function is f(x) = -9x^2 + 18x + 19